# Optimization of Prompt Learning via Multi-Knowledge Representation for Vision-Language Models

### v2

@@ 3.1 Text Encoder Input
Multi-Knowledge  We have defined three different types of Multi-Knowledge: Visual Knowledge (VK), Non-Visual Knowledge (NVK), and Panoramic Knowledge (PK: VK and NVK). The first one is visual knowledge (VK), which includes captions describing the image or its category. Secondly, and of greater importance, we introduce non-visual knowledge (NVK) at a more abstract level beyond purely visual aspects. Thirdly, we can combine multilevel descriptions, such as both VK and NVK, into a more comprehensive description called panoramic knowledge (PK). More instances of Multi-Knowledge can be found in Appendix. The differences between these different types of Multi-Knowledge lie in providing varied perspectives of knowledge description for the same image, beyond solely visual aspects. We generate the Multi-Knowledge using the large language model GPT-4 by constructing different prompts. As shown in Figure 2, the specific method for generating NVK is illustrated. When CLASS = airplane is input into the Prompt we constructed, we obtain the response from GPT-4: Capability to fly carrying passengers or cargo long distances. This represents the Multi-Knowledge described from a non-visual perspective, which is NVK. We utilize GPT-4 to construct Multi Knowledge for downstream tasks as input to the Text Encoder for training. In Appendix, we provide more detailed prompts for generating Multi-Knowledge.
@@
Soft Prompt  In CoKnow, we replace text with a set of continuous vectors as learnable templates. Let $M$ be learnable vectors indexed as $\{v_{1},v_{2},...,v_{M}\}$, where each vector corresponds to an embedding vector for a word. Let $t_{i}$ represent the learnable prompt constructed for the $i$-th class, such that $t_{i}=\{v_{1},v_{2},...,v_{M},c_{i}\}$, where $c_{i}$ represents the embedding vector for the class name of the $i$-th category. Similar to Hand-crafted Templates, each class name shares the same learnable prompt. As shown in Figure 2, the learnable context refers to Soft Prompt. We construct such Soft Prompts to learn effective templates adapted to downstream tasks.
3.2 Training of CoKnow
@@
In this section, we have introduced the complete training framework of CoKnow. The entire training framework is illustrated in Figure 2, where we have only introduced lightweight semantic knowledge mappers as additional components to complete the entire architecture. In the text encoder, three types of templates are inputted: Learnable Context, Multi-Knowledge, and Hand-Crafted Template. At the output end of the text encoder, three sets of representation vectors are obtained: $T_{0}=\{t_{1}^{0},t_{2}^{0},...,t_{k}^{0}\}$, $T_{1}=\{t_{1}^{1},t_{2}^{1},...,t_{k}^{1}\}$, and $T_{2}=\{t_{1}^{2},t_{2}^{2},...,t_{k}^{2}\}$. Here, $k$ represents the number of categories. After a single image is inputted into the image encoder, it is duplicated into two identical visual representation vectors. These vectors then undergo processing through two semantic knowledge mappers, resulting in output vectors $I_{1}$ and $I_{2}$. The logits are calculated between $I_{1}$ and $T_{1}$, as well as between $I_{2}$ and $T_{2}$. Simultaneously, the vectors $I_{0}$, $I_{1}$, and $I_{2}$ are added together using weight parameters $\beta$ to obtain the vector $I^{\prime}$. Finally, logits are calculated between the vector $I^{\prime}$ and $T_{0}$. The specific formulas are as follows:
$\text{logits}_{1}=\text{logit\_scale}\cdot I_{1}T_{1}^{T},$ (1)
$\text{logits}_{2}=\text{logit\_scale}\cdot I_{2}T_{2}^{T},$ (2)
@@
[…]
where the logit_scale refers to a scaling factor. The $\text{logits}_{1}$, $\text{logits}_{2}$, and logits are computed with the label using cross-entropy loss. Finally, the sum of the losses from these three logits is used to update the parameters. We only train the parameters of two lightweight semantic knowledge mappers $w_{1}$ and $w_{2}$ to optimize the context templates. Our semantic knowledge mappers adopt a three-layer fully connected neural network structure, consisting of an input layer, one hidden layer, and an output layer. We set the dimensionality of the hidden layer to one-fourth of the input dimension and add a ReLU activation function after the input layer.
Table 2: Per-dataset results on the ResNet-50 backbone. We bold the best result for each shot and each dataset. CoKnow consistently produces the best performance across all datasets.
@@
Figure 4: The inference stage of CoKnow. The prediction process of CoKnow involves automatically generating additional semantic knowledge representations when an image is inputted. These representations are then aggregated via $\beta$ weights to produce a new vector for prediction. During inference, our method only requires the input image and the completed training contextual template, without any additional inputs needed.
3.3 Inference of CoKnow
In this section, we have introduced the complete inference framework of CoKnow. As illustrated in Figure 4, our method is very concise during the inference stage, requiring only the input of the image and the learned context template, without any additional input. Assuming there are $K$ categories in total, where $w_{i}$ represents the output of the text encoder when the category is $i$, $x_{0}$ represents the output of the image encoder for the given image $I$, and $x_{1}$ and $x_{2}$ represent the outputs of the semantic knowledge mappers $W$ for the given image $I$, the probability of it belonging to category $i$ is as follows:
@@ 3.3 Inference of CoKnow
4 Experiments
Table 3: Comparison with previous methods. We used ResNet50 as the backbone to compare the average top-1 accuracy across 11 test sets under the condition of few-shot learning with five different sample sizes. Linear Probe refers to the Linear-Probe CLIP method. The best results are highlighted in bold, and the second-best results are underlined.
@@ 4 Experiments
4.1 Experimental Settings
Datasets  We conducted experiments on 11 publicly available datasets, encompassing a wide range of scenes and rich content, including ImageNet [6], Caltech [7], Oxford-Pets [21], Flowers [20], Food101 [3], Stanford Cars [15], FGVCAircraft [17], EuroSAT [11], UCF101 [26], DTD [5], and SUN397 [29]. In strict adherence to the few-shot evaluation protocol outlined in CoOp [36], we utilized 1, 2, 4, 8, and 16 shots for training, respectively, and validated the results on the complete test sets.
Implementation Details  To ensure a fair comparison, we strictly adhered to the CoOp protocol [36], employing ResNet-50 as the backbone architecture for the CLIP image encoder. Additionally, we evaluated the performance using ViT-B/16 as the vision encoder. Our experiments were conducted with a batch size of 32, epoch sizes ranging from 10 to 400, and the $\beta$ value set to 0.6. Learning rates of 0.001 and 0.002 were utilized, alongside the SGD optimizer and a cosine annealing strategy. The training process was executed on 6 NVIDIA GeForce RTX 3090 GPUs. We utilize three types of Multi-Knowledge through GPT4, including visual knowledge (VK), non-visual knowledge (NVK), and panoramic knowledge (PK: VK and NVK).
Table 4: Comparison on Distribution Shift. In the context of distribution shift, we have conducted comparisons with previous works. We employed ViT-B/16 for training on the ImageNet dataset, followed by zero-shot testing on the ImageNet-V2 datasets.
@@ 4.1 Experimental Settings
4.2 Performance
@@
Few-shot Learning  In Table 2, we present the comparison of the average top-1 accuracy rates of CoKnow with CoOp across each dataset. This intuitively detailed display showcases the performance difference between CoKnow and prompt learning methods solely utilizing image representations, with CoOp serving as our baseline. Notably, in every accuracy metric, we have demonstrated superior performance. In Table 3, we have compared our approach with a greater variety of methods than before. Our method surpasses a range of previous approaches from 2 shots to 16 shots. It is worth noting that in the 1 shot scenario, CoKnow outperforms both CoOp and Wise-FT’s 2 shots results. With 4 shots, CoKnow’s results approach those of CoOp with 8 shots and surpass Wise-FT’s results. In the 8 shots scenario, CoKnow’s results are comparable to CoOp’s and even surpass Wise-FT’s results. These experimental results all demonstrate the effectiveness of CoKnow in learning from limited data.
Distribution Shift  We further evaluated the robustness of CoKnow under out-of-distribution (OOD) conditions. We conducted prompt learning on ImageNet and tested on ImageNetV2 [24]. The results are presented in Table 4. The experimental results indicate that the method effectively generalizes from ImageNet to out-of-distribution datasets, showcasing its potential in handling distribution shifts.
Figure 5: Ablation Study on Inputting Different Multi-Knowledge. At the text input end, apart from the Hand-crafted Template, we input different Multi-Knowledge, including NVK, VK, and PK, respectively, as target Multi-Knowledge Representation for learning. The Baseline represents direct prompt learning via CoOp.
@@
4.3 Ablation Study
Table 5: Ablation Study of Mapper. CoKnow-I represents the re-mapping of the original CLIP’s image representations in CoKnow using a semantic knowledge mapper.
@@ 4.3 Ablation Study
Table 6: Ablation Study on Different Classname Positions. End represents the category name located at the end of the learnable vectors, while Middle represents the category name located in the middle of the learnable vectors.
@@
Table 7: Average Prediction Results for Different Context Lengths. This represents the lengths of learnable vectors of different sizes.
@@
Different Multi-Knowledge.  We utilize VK, NVK, and PK, three different types of knowledge descriptions, to further explore the impact of different Multi-Knowledge on prompt learning results under the condition of inputting the same Hand-crafted Template. Setting $\beta$ at 0.8 and optimizing the context length to 16, we detailed the average effects of few-shot learning across 11 datasets in Figure 5. Overall, descriptions of varying Multi-Knowledge were able to further optimize prompt learning, achieving commendable levels of performance. The overall prediction accuracy of PK is the best, indicating that a global representation incorporating knowledge from multiple perspectives is more effective.
@@
Different $\beta$ weight values.  We investigated the impact of different $\beta$ weights on the model’s outcomes by setting the $\beta$ values to 0.4, 0.6, and 0.8, respectively, with PK as the input and the optimization context length set to 16. Figure 6 presents a comparison of the average predictions across 11 datasets under the condition of 16 shots. Simultaneously, the effects of previous methods are directly demonstrated in the graph. Different values of $\beta$ have a certain impact on the results, but they all outperform a series of previous excellent methods. Our experiments offer guidance for adjusting the $\beta$ value.
@@
Ablation Study of Mapper.  The experiment explores the effectiveness of using a semantic knowledge mapper to map the image representations of the original CLIP (denoted as $I_{0}$ in Figure 2), as well as investigates the importance of the original CLIP’s image representations for Prompt Learning. The $\beta$ value is set to 0.6, the input is set to PK, and the optimized context length is set to 16. We incorporate a semantic knowledge mapper before $I_{0}$ in Figure 2, and the new structure is referred to as CoKnow-I. As shown in Table 5, it can be observed that when the training sample size is 1 shot, CoKnow-I achieves very low Top-1 accuracy across 11 datasets. When the training sample size is increased to 16 shots, the Top-1 accuracy across these 11 datasets remains lower than the predicted results of CoKnow. The experimental results indicate that the image representations of the original CLIP have a significant impact on Prompt Learning, especially evident when the training sample size is 1-shot. Additionally, it is not advisable to use a semantic knowledge mapper to map the image representations of the original CLIP.

### v3

@@ III-A2 Multi-Knowledge
We have defined three different types of Multi-Knowledge: visual knowledge (VK), nonvisual knowledge (NVK), and panoramic knwoledge (PK: VK and NVK). The first is visual knowledge (VK), which includes captions describing the image or its category. Second, and of greater importance, we introduce nonvisual knowledge (NVK) at a more abstract level beyond purely visual aspects. Third, we can combine multilevel descriptions, such as both VK and NVK, into a more comprehensive description called panoramic knowledge (PK). We generate the Multi-Knowledge via the large language model GPT-4 by constructing different prompts. As shown in Figure 4, the specific method for generating the NVK is illustrated. When CLASS = airplane is input into the prompt we constructed, we obtain the response from GPT-4: Capability to fly carrying passengers or cargo long distances. This represents the Multi-Knowledge described from a nonvisual perspective, which is NVK. We utilize GPT-4 to construct Multi-Knowledge for downstream tasks as input to the text encoder for training. In Figure 7, we illustrate the pipeline for generating Multi-Knowledge via GPT-4. We present a specific instance to generate a PK utilizing GPT-4. Upon providing a category face to GPT-4 as input, the comprehensive prompt yields the following response: Area with eyes, nose, mouth; center of emotional expression.
III-A3 Soft Prompt
@@
In CoKnow, we replace text with a set of continuous vectors as learnable templates. Let $M$ be learnable vectors indexed as $\{v_{1},v_{2},...,v_{M}\}$, where each vector corresponds to an embedding vector for a word. Let $t_{i}$ represent the learnable prompt constructed for the $i$-th class such that $t_{i}=\{v_{1},v_{2},...,v_{M},c_{i}\}$, where $c_{i}$ represents the embedding vector for the class name of the $i$-th category. Like hand-crafted templates do, each class name shares the same learnable prompt. As shown in Figure 4, the learnable context refers to the soft prompt. We construct such soft prompts to learn effective templates adapted to downstream tasks.
Figure 7: The specific process of generating Multi-Knowledge. By providing the given category name along with a fixed prompt, we input these names into GPT-4 to generate a response.
III-B Training of CoKnow
@@
In this section, we introduce the complete training framework of CoKnow. The entire training framework is illustrated in Figure 4, where we introduce only lightweight semantic knowledge mappers as additional components to complete the entire architecture. In the text encoder, three types of templates are input: learnable context, Multi-Knowledge, and hand-crafted template. At the output end of the text encoder, three sets of representation vectors are obtained: $T_{0}=\{t_{1}^{0},t_{2}^{0},...,t_{k}^{0}\}$, $T_{1}=\{t_{1}^{1},t_{2}^{1},...,t_{k}^{1}\}$, and $T_{2}=\{t_{1}^{2},t_{2}^{2},...,t_{k}^{2}\}$. Here, $k$ represents the number of categories. After a single image is input into the image encoder, it is duplicated into two identical visual representation vectors. These vectors then undergo processing through two semantic knowledge mappers, resulting in output vectors $I_{1}$ and $I_{2}$. The logits are calculated between $I_{1}$ and $T_{1}$, as well as between $I_{2}$ and $T_{2}$. Simultaneously, the vectors $I_{0}$, $I_{1}$, and $I_{2}$ are added together via weight parameters $\beta$ to obtain the vector $I^{\prime}$. Finally, logits are calculated between the vectors $I^{\prime}$ and $T_{0}$. The specific formulas are as follows:
$\text{logits}_{1}=\lambda\cdot I_{1}T_{1}^{T},$ (1)
$\text{logits}_{2}=\lambda\cdot I_{2}T_{2}^{T},$ (2)
@@
$\text{logits}=\lambda\cdot I^{\prime}T_{0}^{T},$ (4)
where $\lambda$ refers to a scaling factor. The $\text{logits}_{1}$, $\text{logits}_{2}$, and logits are computed with the label via cross-entropy loss. Finally, the sum of the losses from these three logits is used to update the parameters. We only train the parameters of two lightweight semantic knowledge mappers $w_{1}$ and $w_{2}$ to optimize the context templates. Our semantic knowledge mappers adopt a three-layer fully connected neural network structure consisting of an input layer, one hidden layer, and an output layer. We set the dimensionality of the hidden layer to one-fourth of the input dimension and add a ReLU activation function after the input layer. Specifically, we set the input layer dimension of the semantic knowledge mappers to correspond with the hidden size of the visual encoder (e.g., 1024). The intermediate layer dimension is set to one-fourth of the input layer dimension (e.g., 256), and the output layer is remapped to the same dimension as the input layer (1024). Additionally, we add ReLU activation functions between the input and intermediate layers, as well as between the intermediate and output layers.
Figure 8: The inference stage of CoKnow. The CoKnow prediction process involves automatically generating additional semantic knowledge representations when an image is input. These representations are then aggregated via $\beta$ weights to produce a new vector for prediction. During inference, our method only requires the input image and the completed training contextual template, without any additional inputs needed.
TABLE II: Datasets statistics.
Dataset Classes Train Val Test
ImageNet 1,000 1.28M N/A 50,000
Caltech101 100 4,128 1,649 2,465
OxfordPets 37 2,944 736 3,669
StanfordCars 196 6,509 1,635 8,041
Flowers102 102 4,093 1,633 2,463
Food101 101 50,500 20,200 30,300
FGVCAircraft 100 3,334 3,333 3,333
SUN397 397 15,880 3,970 19,850
DTD 47 2,820 1,128 1,692
EuroSAT 10 13,500 5,400 8,100
UCF101 101 7,639 1,898 3,783
CIFAR-10 10 50,000 N/A 10,000
CIFAR-100 100 50,000 N/A 10,000
STL-10 10 5,000 N/A 8,000
TABLE III: Per-dataset results on the ResNet-50 backbone. We bold the best result for each shot and each dataset. CoKnow consistently produces the best performance across all datasets.
@@
III-C Inference of CoKnow
In this section, we introduce the complete inference framework of CoKnow. As illustrated in Figure 8, our method is very concise during the inference stage, requiring only the input of the image and the learned context template, without any additional input. Assuming that there are $K$ categories in total, where $w_{i}$ represents the output of the text encoder when the category is $i$, $x_{0}$ represents the output of the image encoder for the given image $I$, and $x_{1}$ and $x_{2}$ represent the outputs of the semantic knowledge mappers $W$ for the given image $I$; the probability of it belonging to category $i$ is as follows:
@@ III-C Inference of CoKnow
IV Experiments
TABLE IV: Comparison with previous methods. We used ResNet50 or ViT-B/16 as the backbone to compare the average top-1 accuracy across 11 test sets under the condition of few-shot learning with five different sample sizes. The best results are highlighted in bold, and the second-best results are underlined.
@@ IV Experiments
ViT-B/16
CoCoOp CVPR’22 69.10 70.38 72.32 76.20 78.43
Tip-Adapter ECCV’22 69.81 71.56 74.18 75.17 77.39
Cross-Modal WiSE-FT CVPR’23 70.08 71.13 72.97 76.96 79.45
KgCoOp CVPR’23 69.30 70.68 72.51 76.55 78.71
CoKnow - 69.44 72.67 74.93 77.80 80.78
TABLE V: Comparison on the distribution shift. In the context of distribution shift, we have conducted comparisons with previous works. We employed ViT-B/16 for training on the ImageNet dataset, followed by zero-shot testing on the ImageNet-V2 datasets.
@@
IV-A Experimental Settings
IV-A1 Datasets
We conducted experiments on 11 publicly available datasets, encompassing a wide range of scenes and rich content, including ImageNet [36], Caltech [35], Oxford-Pets [42], Flowers [41], Food101 [40], Stanford Cars [43], FGVCAircraft [39], EuroSAT [38], UCF101 [45], DTD [37], and SUN397 [44]. In Table 6, we present the statistical information of the datasets. In strict adherence to the few-shot evaluation protocol outlined in CoOp [13], we utilized 1, 2, 4, 8, and 16 shots for training and validated the results on the complete test sets.
IV-A2 Implementation Details
To ensure a fair comparison, we strictly adhered to the CoOp protocol [13], employing ResNet-50 or ViT-B/16 as the backbone architecture for the CLIP image encoder. Additionally, we evaluated the performance using ViT-B/16 as the vision encoder. The language model used maintains the same temperature parameter across each dataset, set to the default value. Our experiments were conducted with a batch size of 32 and epoch sizes ranging from 10 to 400. Learning rates of 0.001 and 0.002 were utilized alongside the SGD optimizer and a cosine annealing strategy. The training process was executed on 6 NVIDIA GeForce RTX 3090 GPUs. We utilize three types of Multi-Knowledge through GPT-4 (GPT-4 refers to the large language model endpoint ”gpt-4” developed by OpenAI.) These include visual knowledge (VK), nonvisual knowledge (NVK), and panoramic knowledge (PK: VK and NVK). In the experiments of Few-shot Learning and Distribution Shift, we set the $\beta$ value at 0.6, which was determined to yield the optimal experimental results. In the ablation studies, we accordingly explored the impact of varying $\beta$ values on the experiments, and the specific $\beta$ values employed can be found in the corresponding ablation study sections.
IV-B Performance
@@
Figure 9: Ablation study on inputting different Multi-Knowledge. At the text input end, in addition to the hand-crafted template, we input different muultiknowledge, including NVK, VK, and PK, as target multi-knowledge representations for learning. The Baseline represents direct prompt learning via CoOp.
TABLE VI: Multi-Knowledge Ablation Experiment ($\beta=0.6$).
Multi-knowledge Number of Shots
NVK VK PK 1 2 4 8 16
 – 59.59 62.32 66.77 69.89 73.42
 ✓ 63.10 66.12 69.73 72.75 75.68
63.01 66.03 69.35 72.76 75.65
 ✓ 63.38 66.49 69.87 73.04 76.09
IV-B1 Few-shot Learning
In Table III, we present a comparison of the average top-1 accuracy rates of CoKnow with CoOp across each dataset. This intuitively detailed display highlights the performance difference between CoKnow and prompt learning methods solely utilizing image representations, with CoOp serving as our baseline. Notably, for every accuracy metric, we have demonstrated superior performance. In Table IV, we compared our approach with a greater variety of methods than before. Our method surpasses a range of previous approaches from 2 shots to 16 shots. Notably, in the 1-shot scenario, CoKnow outperforms both CoOp and Wise-FT in terms of the 2 shot results. With 4 shots, CoKnow’s results approach those of CoOp with 8 shots and surpass the results of Wise-FT. In the 8-shot scenario, the CoKnow results are comparable to those of CoOp and even surpass those of Wise-FT. These experimental results demonstrate the effectiveness of CoKnow in learning from limited data.
IV-B2 Distribution Shift
We further evaluated the robustness of CoKnow under out-of-distribution (OOD) conditions. We conducted prompt learning on ImageNet and tested it on ImageNetV2 [46]. The results are presented in Table V. The experimental results indicate that the method effectively generalizes from ImageNet to out-of-distribution datasets, highlighting its potential in handling distribution shifts.
IV-B3 Inference Overhead
We present the costs of training and inference for CoKnow on datasets such as Caltech101, Pets, and Cars in Table VII. The table presents the results when the batch size is 32 and the learnable prompt length is 16. The lightweight framework of CoKnow effectively reduces the inference overhead, making it easier to apply to downstream tasks. Specifically, the semantic knowledge mapper in the CoKnow architecture allows us to avoid loading large amounts of Multi-Knowledge during the inference phase, thereby effectively reducing the inference overhead. For example, the memory usage on the StanfordCars dataset decreases from 4002 MB to 1434 MB.
TABLE VII: Overhead during training and inference.
Datasets Training (MB) Inference (MB) $\downarrow$
OxfordPets 1370 932
Caltech101 2488 988
StanfordCars 4002 1434
@@ IV-B3 Inference Overhead
IV-C Ablation Study
IV-C1 Different Multi-Knowledge
We utilize VK, NVK, and PK, three different types of knowledge descriptions, to further explore the impact of different Multi-Knowledge on prompt learning results when the same hand-crafted template is input. Setting $\beta$ at 0.8 and optimizing the context length to 16, we detail the average effects of few-shot learning across 11 datasets in Figure 9. Overall, descriptions of varying Multi-Knowledge were able to further optimize prompt learning, achieving commendable levels of performance. The overall prediction accuracy of PK is the best, indicating that a global representation incorporating knowledge from multiple perspectives is more effective. We present the impact of different Multi-Knowledge scenarios with $\beta$ set to 0.6 in Table VI for further demonstration.
TABLE VIII: Ablation Study of Mapper. CoKnow-I represents the re-mapping of the original CLIP’s image representations in CoKnow using a semantic knowledge mapper.
@@ IV-C1 Different Multi-Knowledge
TABLE IX: Ablation Study on Different Classname Positions. End represents the category name located at the end of the learnable vectors, whereas Middle represents the category name located in the middle of the learnable vectors.
@@
TABLE X: Average Prediction Results for Different Context Lengths. This represents the lengths of learnable vectors of different sizes.
@@
IV-C2 Different $\beta$ weight values.
@@
We investigated the impact of different $\beta$ weights on the model’s outcomes by setting the $\beta$ values to 0.4, 0.6, and 0.8, with PK as the input and the optimization context length set to 16. Figure 10 presents a comparison of the average predictions across 11 datasets under the condition of 16 shots. Simultaneously, the effects of previous methods are directly demonstrated in the graph. Different values of $\beta$ have a certain impact on the results, but they all outperform a series of previous excellent methods. Our experiments offer guidance for adjusting the $\beta$ value.
IV-C3 Ablation Study of Mapper
@@
This experiment explores the effectiveness of using a semantic knowledge mapper to map the image representations of the original CLIP (denoted as $I_{0}$ in Figure 4) and investigates the importance of the original CLIP’s image representations for prompt learning. The $\beta$ value is set to 0.6, the input is set to PK, and the optimized context length is set to 16. We incorporate a semantic knowledge mapper before $I_{0}$ in Figure 4, and the new structure is referred to as CoKnow-I. As shown in Table VIII, when the training sample size is 1 shot, CoKnow-I achieves very low Top-1 accuracy across 11 datasets. When the training sample size is increased to 16 shots, the Top-1 accuracy across these 11 datasets remains lower than the predicted results of CoKnow. The experimental results indicate that the image representations of the original CLIP have a significant effect on prompt learning, especially when the training sample size is 1 shot. Additionally, it is not advisable to use a semantic knowledge mapper to map the image representations of the original CLIP.
IV-C4 Position of Classname Varies